# End-to-End Speech-to-Text Translation: A Survey

### v1

@@ 6.1 Cascade vs End-to-End Models
As argued and presented through comprehensive experiments by (Bentivogli et al., 2021), the performance gaps between cascade and E2E ST models are bridged. However, as shown by (Agrawal et al., 2023) in a recent IWSLT 2023 subtitling generation task, the performance of cascade models is far superior to E2E models for offline ST tasks evaluated on all metrics. Furthermore, as far as our understanding, no thorough assessment has been done for low-resource languages that use E2E and cascade models. It may be interesting to compare E2E and cascade ST models on various ST datasets.
@@ 6.2 ST on Code-Mix data
We find that there is only one study on the ST model that uses code-mix data as an input. A code-mix data has problems, such as different lexicons, syntax, and scarcity of labeled data. Therefore, it will be interesting to (a) create Code-Mix ST datasets incorporating more languages, (b) see how the existing ST models perform on code-mix ST data?, and (c) Can pre-training in many languages assist in tackling the code-mixing issue.
@@ 6.4 Discrepancy between Automatic and Human Evaluation
There may be discrepancies and disagreements among various metrics used to report ST task results. They do not match the mean option score (MOS) provided by human evaluators (Agrawal et al., 2023). For example, if a system evaluates the BLUE score between a ground truth sentence “Police shot the culprit with a gun” and hypothesis sentence “Police use a gun to shot the culprit”, it is 0! However, both sentences above might be deemed appropriate translations of an utterance semantically by an ST system. Such an argument is supported by dubbing artists who often change the voice of the sentence to simplify it or make it more pleasing.  [^17] In the movie “Pirates of the Caribbean”, Jack Sparrow asks Bloom how long he can go for the girl. The original answer from Bloom is “I can die for her!”. Whereas Hindi dubbing is “Till the dying breadth”
Therefore, we call for the attention of researchers to develop and use metrics that match human evaluations semantically. An approach could be to subject the ground truth and hypothesis sentences under semantic textual similarity tasks and score them accordingly.
@@ 6.8 Multilingual and Simultaneous ST
Multilingual ST has gained momentum recently due to its importance in the real world. For example, a single speech must be broadcast to multilingual communities (e.g., a conference is attended by a diverse group of people). It can be one-to-many, many-to-one, and many-to-many languages ST. Our literature survey shows that only a few works exist to target one-to-many and many-to-one, whereas many-to-many is few (Inaguma et al., 2019). Besides, there is an opportunity to explore simultaneous multilingual ST, which is the most practical setting.
@@ 6.9 Low-resource ST Datasets and Models
Most existing works have focused on building ST models and datasets for high-resource languages. As we know, the success of ST models relies on the parallel speech-text corpus; building ST datasets for low-resource languages requires immediate attention. Further, a few works, such as (Bansal et al., 2018), have reported ST task results on the Mboshi-French pair; however, the BLUE score is poor. Therefore, building models that transfer learning from language pairs with high to low resources is warranted.

### v2

@@ 10.1 Cascade vs End-to-End Models
As argued and presented through comprehensive experiments by (Bentivogli et al., 2021), the performance gaps between cascade and E2E ST models are bridged. However, as shown by (Agrawal et al., 2023) in a recent IWSLT 2023 subtitling generation task, the performance of cascade models is far superior to E2E models for offline ST tasks evaluated on all metrics. Furthermore, as far as our understanding, no thorough assessment has been done for low-resource languages that use E2E and cascade models. It may be interesting to compare E2E and cascade ST models on various ST datasets to assert the claims in the literature.
@@ 10.2 ST on Code-Mix data
We find that there exists limited study on the ST model that uses code-mix data as an input. A code-mix data has problems, such as different lexicons, syntax, and scarcity of labeled data. Therefore, it will be interesting to (a) create Code-Mix ST datasets incorporating more languages, (b) see how the existing ST models perform on code-mix ST data?, and (c) Can pre-training in many languages assist in tackling the code-mixing issue?
@@ 10.4 Discrepancy between Automatic and Human Evaluation
There may be discrepancies and disagreements among various metrics used to report ST task results. They do not match the mean option score (MOS) provided by human evaluators (Agrawal et al., 2023). For example, if a system evaluates the BLEU score between a ground truth sentence “Police shot the culprit with a gun” and hypothesis sentence “Police use a gun to shot the culprit”, it is 0! However, both sentences above might be deemed appropriate translations of an utterance semantically by an ST system. Such an argument is supported by dubbing artists who often change the voice of the sentence to simplify it or make it more pleasing.  [^22] In the movie “Pirates of the Caribbean”, Jack Sparrow asks Bloom how long he can go for the girl. The original answer from Bloom is “I can die for her!”. Whereas Hindi dubbing is “Till the dying breadth”
As highlighted in (Marie et al., 2021), the BLEU score is being reported by more than 99% of MT papers without accounting for statistical significance testing or human evaluation. Our survey of ST papers indicates the same trend being followed. Therefore, we call for the attention of researchers to develop and use metrics that match human evaluations semantically. An approach could be to subject the ground truth and hypothesis sentences under semantic textual similarity tasks and score them accordingly.
@@ 10.8 Multilingual and Simultaneous ST
Multilingual ST has gained momentum recently due to its importance in the real world. For example, a single speech must be broadcast to multilingual communities (e.g., a conference is attended by a diverse group of people). It can be one-to-many, many-to-one, and many-to-many languages ST. Our literature survey shows that only a few works exist in this space. Besides, there is an opportunity to explore simultaneous multilingual ST, which is the most practical setting.
@@ 10.9 Low-resource ST Datasets and Models
Most existing works have focused on building ST models and datasets for high-resource languages. As we know, the success of ST models relies on the parallel speech-text corpus; building ST datasets for low-resource languages requires more attention. Further, a few works, such as (Bansal et al., 2019), have reported ST task results on the Mboshi-French pair; however, the BLEU score is poor. Therefore, building models that transfer learning from language pairs with high to low resources is warranted.
10.10 LLMs for ST tasks
In the last few years, large language models (LLMs) have emerged as a promising solution to many NLP tasks including ST. LLMs show in-context learning (ICL) when trained over a massive amount of data. This process unlocks their hidden emergent abilities (Wei et al., 2022) and enables them for few-shot and zero-shot learning capability via prompting. There exist a few works (Zhang et al., 2023b; Wu et al., 2023; Huang et al., 2023) (see (Gaido et al., 2024) for a comparative discussion) which explore LLMs for ST task. Concretely, all of these models leverage a speech foundation model (SLM) followed by length adapter, modality adaptation, mixing the two modalities, and then LLMs for generating the output. GenTranslate (Hu et al., 2024) builds upon the Seamless4MT by integrating an LLM on top and performing $N$-best hypothesis tuning. Initial results are plausible. However, it remains to see how various components affect the downstream task performance, what is the best strategy for prompt design, and how to pre-train/fine-tune them in a parameter-efficient way for ST tasks. Further, the use of LLMs for SimulMT has been recently proposed (Agostinelli et al., 2023) and it remains to see how to adapt SimulMT to SimulST.
10.11 Really long Context Modelling
As mentioned in the streaming section, SST models need to handle long input sequences. Current speech encoders lack infinite context modeling capability due to their quadratic complexity of self-attention. There have been recent improvements to handle the problem of infinite context. For example, Mamba (Zhang et al., 2024a), Infini-attention (Munkhdalai et al., 2024), and TransforerFAM (Hwang et al., 2024) show some promising results in long context modeling. These models may be explored for SST task as well.